If the dissociation constant of HCN is 4.8 X 10-10 mol dm-3, then the pH of 0.25 mol dm-3 NaCN solution is



Grade:

1 Answers

Askiitians Expert Sahil Arora - IITD
13 years ago

Dear Nehal

As NaCN is a salt of strong base and weak acid

So pH =  1/2pKw +  1/2pKa +  1/2log C

= 7 + 1/2pKa +  1/2log C.

 

 

 

 

Put Ka=4.8 X 10-10 mol dm-3 and C=0.25 mol/dm3.

we get pH=8.55

for more details read:-

http://www.askiitians.com/iit-jee-chemistry/physical-chemistry/salt-of-weak-acid-and-strong-base.aspx

Please feel free to post as many doubts on our discussion forum as you can. If you find any question 
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Nehal !!!




Regards,

Askiitians Experts

Sahil Arora